OpenStudy (anonymous):

a rectangle garden has dimensions 25 ft by 20 ft. A gravel path of equal width is to be built around the garden. How wide can the path be if there is enough gravel for 550 sq ft.

OpenStudy (anonymous):

w=22

OpenStudy (anonymous):

let the horizontal dimension be 25 and the vertical be 20. then draw your path around the rectangle. There will be two ears on the left and right of area 20*w. There are two ears top and bottom area 25*w. Then there are 4 corners area w^2. In total: 40w+50w+4w^2=550 4w^2+90w=550 w=5 ft.

OpenStudy (anonymous):

oh. i see thanks.

OpenStudy (anonymous):

it is just a matter of using A=l*w then substitute.

OpenStudy (anonymous):

what do you mean?